OpenStudy (anonymous):

(2x2−18)÷(18−6x) what is the numerator?

OpenStudy (anonymous):

***(2x^2 - 18)

jimthompson5910 (jim_thompson5910):

\[\large \frac{2x^2-18}{18-6x}\] \[\large \frac{2(x^2-9)}{-6(x-3)}\] \[\large \frac{2(x-3)(x+3)}{-6(x-3)}\] \[\large \frac{2(x+3)}{-6}\] \[\large -\frac{x+3}{3}\] So \[\large \frac{2x^2-18}{18-6x}=-\frac{x+3}{3}\]

OpenStudy (aroub):

2(x^2-9)/6(x-3) 2(x-3)(x+3)/6(x-3) x+3/3(x-3) the numerator is x+3